# Cross-Image Contrastive Decoding: Precise, Lossless Suppression of Language Priors in Large Vision-Language Models

### v3

@@ Abstract
Language priors are a major cause of hallucinations in Large Vision-Language Models (LVLMs), often leading to text that is linguistically plausible but visually inconsistent. Recent work explores contrastive decoding as a training-free solution, but these methods typically construct negative contexts from the original image, resulting in visual information loss and distorted distribution. Motivated by the observation that language priors stem from the LLM backbone and remain consistent across images, we propose Cross-Images Contrastive Decoding (CICD), a simple yet effective training-free method that uses different images to construct negative contexts. We further analyze the cross-image behavior of language priors and introduce a distinction between essential priors (supporting fluency) and detrimental priors (causing hallucinations). By selectively preserving essential priors and suppressing detrimental ones, our method reduces hallucinations while maintaining coherent and fluent language generation. Experiments on 4 benchmarks and 6 LVLMs across three model families confirm the effectiveness and generalizability of CICD, especially in image captioning, where language priors are particularly pronounced. Code will be released once accepted.
@@
Jianfei Zhao1,2, Feng Zhang1, Xin Sun1, Chong Feng1,3,thanks: Corresponding Autor: fengchong@bit.edu.cn , 1School of Computer Science and Technology, Beijing Institute of Technology, Beijing, China, 2Zhongguancun Academy, Beijing, China, 3Southeast Academy of Information Technology, Beijing Institute of Technology, Fujian, China,
@@ 1 Introduction
Figure 1: A case of language prior. Adding the same prefix when captioning another image will lead to the same language priors. The model used is LLaVA-1.5.
@@
Recent advances in Large Language Models (LLMs) Brown et al. (2020); Touvron et al. (2023); Chiang et al. (2023); Bai et al. (2023a) have opened promising pathways toward Artificial General Intelligence (AGI). Building on these developments, numerous studies Liu et al. (2023); Dai et al. (2023); Bai et al. (2023b, 2025) have extended LLMs by integrating visual inputs, giving rise to Large Vision-Language Models (LVLMs) capable of understanding visual content. They are capable of addressing a broad spectrum of vision-language tasks, such as image captioning Jing et al. (2024); Rohrbach et al. (2018); Ye et al. (2025); Chen et al. (2025), visual question answering (VQA) Lovenia et al. (2024); Li et al. (2023b); Yu et al. (2024), and visual grounding Rasheed et al. (2024); You et al. (2024); Lu et al. (2024).
Despite impressive capabilities, LVLMs remain susceptible to hallucinations Li et al. (2023c); Wang et al. (2023), often generating text that contradicts the visual input. These hallucinations stem from various sources, with language priors Chen et al. (2025); Leng et al. (2024); Wang et al. (2024c); Zhu et al. (2024b) being the most fundamental underlying cause. Specifically, since LVLMs are built upon LLMs, they inherently inherit the linguistic knowledge encoded in the LLMs. As a result, their outputs are influenced not only by visual evidence but also by pre-learned language patterns. This often leads to a tendency to rely on memorized linguistic associations when responding to user queries. For instance, when describing an office scene, an LVLM may erroneously mention a "computer" even if no such object is present in the image.
Figure 2: The overlap ratio of image-related words between responses generated from negative contexts and those from regular inputs. Both VCD and ICD exhibit over 30% overlap, indicating that their negative contexts retain a considerable amount of essential visual information. In contrast, our method results in a nearly negligible overlap. The detailed analysis methodology can be found in Appendix. D.
Methods for mitigating language priors in LVLMs can generally be classified into two categories. The first involves constructing specialized training datasets to reduce the model’s dependence on linguistic biases Zhibo et al. (2023); Wu et al. (2022); Chen et al. (2025), though such approaches often incur substantial training costs. Another line of work explores plug-and-play, training-free methods, which are also the focus of our study. Among them, many recent studies adopt contrastive decoding Li et al. (2023a), which constructs negative contexts to reveal and suppress language priors. Some approaches directly perturb the input image Leng et al. (2024); Wang et al. (2024c) to alter the visual context, while others enhance the visual features extracted from the original image Zhu et al. (2024b); Zhang et al. (2025). These methods typically compute the difference between the logit distributions obtained from the original and perturbed (i.e., negative) contexts to reduce the model’s reliance on language priors. Overall, following the paradigm introduced in VCD Leng et al. (2024), these methods routinely perturb the original inputs to generate negative contexts. However, this prevailing design choice raises a critical question: Is perturbing the original context to construct negative contexts truly the most effective strategy for inducing and mitigating language priors?
Figure 3: Essential priors (left) and detrimental priors (right). We randomly select 500 images from the MSCOCO dataset and visualize LLaVA-1.5’s output by the t-SNE algorithm van der Maaten and Hinton (2008). w/ Image represents regular outputs. w/o Image represents the outputs after removing the input image. w/ Neg. Image represents replacing the input image by the most dissimilar image in the dataset retrieved by CLIP Radford et al. (2021).
We argue that the answer is unequivocally negative. Constructing negative contexts based on the original inputs can lead to two main issues: (1) Visual information loss, where the negative context retains partial visual information, which may be degraded during contrastive decoding. As shown in Fig. 2, compared to regular responses, those generated from negative contexts retain over 30% of the essential visual information, which is subsequently weakened during the contrastive decoding process. (2) Distorted distribution, where perturbing the model’s input to induce priors may disrupt its internal representations, ultimately resulting in a distorted logit distribution. This analysis raises an important question: How should negative contexts be constructed to avoid the above issue?
Figure 4: A case of JS divergence between the logit distributions of two images.
In fact, simply taking a different image to construct negative context is sufficient. This insight is based on the thought that the language priors guiding step-wise token generation in LVLMs originate from the LLM backbone and are largely independent of the input image. As shown in Fig.1, LVLM produces highly consistent outputs when given completely different images but the same prefix, which strongly supports our hypothesis. Building on this, we propose Cross-Images Contrastive Decoding (CICD), a simple yet effective method for suppressing language priors in LVLMs. Instead of perturbing the original inputs, CICD constructs the negative context using a different image that shares the same language priors. Our method avoids the two key issues present in previous approaches. Since the overlap between the original and the new image is limited, our approach avoids introducing critical visual information into the negative context. As shown in Fig. 2, only a negligible amount (approximately 0.11%) of essential visual information is retained in our negative context. Moreover, it enables the model to reveal intrinsic priors, maintaining the model’s natural inference process without injecting external bias. The formal analysis of our method is presented in Sec. 3.2.
Focusing on language priors, we further explore whether all such priors should be strictly eliminated. The answer is intuitively negative. Language priors can be divided into two categories: (1) Essential Priors, which are necessary for maintaining textual fluency, and (2) Detrimental Priors, which introduce information unrelated to the visual input. Clearly, only the detrimental priors should be removed. This raises a critical question: How can we accurately distinguish between the two and selectively eliminate only the detrimental ones? As shown in Fig.3, essential priors exhibit high consistency across different images, whereas detrimental priors show substantial divergence when generating image-related content. Based on this observation, we propose to distinguish essential from detrimental priors using cross-image consistency, as illustrated in Fig.4. Specifically, we compute the Jensen–Shannon (JS) divergence between the logit distributions of different visual contexts, and apply a stable threshold to identify essential priors that should be retained.
Extensive experiments on 4 widely used benchmarks and 6 LVLMs across three model families demonstrate the effectiveness and generalizability of our method. By suppressing language priors, CICD consistently reduces hallucinations and achieves state-of-the-art performance across all benchmarks. The improvement is particularly pronounced in the image captioning task, which tends to suffer more from language priors due to its longer textual outputs.
In summary, our main contributions are as follows:
  • We analyze existing contrastive decoding methods in depth and reveal their limitations stemming from reliance on the original image. To address this, we verify the consistency of language priors across images and propose a stable approach to distinguish essential priors from detrimental ones.
  • Based on this analysis, we propose CICD (Cross-Images Contrastive Decoding), which leverages retrieved images to induce language priors while selectively preserving essential priors during the mitigation process.
@@
  • Extensive experiments on multiple benchmarks and LVLMs demonstrate the superiority of our method in alleviating language priors, particularly in the image captioning task.
@@ 2 Related Work
Due to space limitations, we provide a brief overview here. A more detailed discussion of related work can be found in Appendix.A.
@@ 2.1 Hallucination in LVLMs
Hallucinations in LVLMs primarily concern cross-modal consistency Li et al. (2023c); Wang et al. (2023), which requires that the generated content remains faithful to the visual input. Among the various causes of hallucination, language priors Zhibo et al. (2023); Wu et al. (2022) are the most fundamental. These priors originate from the LLM’s pretraining corpus and guide generation based on learned language patterns, which can produce fluent yet visually inconsistent outputs. Several approaches Chen et al. (2025); Liu et al. (2024a); Gunjal et al. (2024) attempt to mitigate language priors by constructing specialized training datasets, but these methods often incur significant additional training costs. Alternatively, recent training-free methods have explored contrastive decoding Li et al. (2023a); Chuang et al. (2024), typically constructing negative contexts based on the original image Leng et al. (2024); Wang et al. (2024c); Zhu et al. (2024b); Zhang et al. (2025) in an attempt to elicit and suppress language priors.
In this work, we identify critical limitations of these approaches that stem from their reliance on modifying the original image. Specifically, such modifications may lead to visual information loss and distorted distribution. We further conduct an in-depth analysis of the cross-image consistency of language priors and, based on this insight, propose CICD. Additionally, we introduce a method for distinguishing between essential priors and detrimental priors, enabling fine-grained and precise mitigation of language priors in LVLMs.
3 Preliminary
@@
Figure 5: The diagram of our method. We input the target image and a distinct image into the LVLM separately. During the step-by-step generation of the response, we compute the Jensen-Shannon (JS) divergence between the two logit distributions at each step to distinguish essential priors from detrimental ones. For detrimental priors, we apply contrastive decoding to suppress harmful language biases, while regular decoding is used for essential priors.
3.1 Contrastive Decoding
LVLMs take as input a concatenation of vision tokens from a vision encoder and an encoded text prompt. After forward propagation, the model projects the output into a logit distribution to compute conditional probabilities and generate responses in an autoregressive manner:
$y_{t}\sim p_{\theta}(y_{t}\mid v,x,y_{<t})$ (1)
$=$ $\ \mathrm{softmax}\left[\mathrm{logit}_{\theta}(y_{t}\mid v,x,y_{<t})\right]$
Here, $v$ and $x$ represent the visual and textual inputs, respectively. $\mathrm{logit}_{\theta}$ represents the logit distribution formulated by the model with parameters $\theta$.
Since LVLMs adopt pretrained LLMs as their backbone, they inherently inherit language priors independent of the visual input. As a result, tokens ranked higher in the output distribution $\mathrm{logit}_{\theta}$ may appear semantically plausible but can be objectively inconsistent with the visual evidence. This constitutes a fundamental cause of hallucinations. A commonly used approach to mitigate such priors is contrastive decoding Li et al. (2023a); Leng et al. (2024), which modifies the original image to construct a native visual context that elicits language priors. These priors are then suppressed by subtracting the logit distribution derived from the native context from that of the original context.
$\mathrm{logit}_{\mathrm{CD}}(y_{t}\mid v,x,y_{<t})$ (2)
$=$ $(1+\alpha)\,\mathrm{logit}_{\theta}(y_{t}\mid v,x,y_{<t})-\alpha\,\mathrm{logit}_{N}$
Here, $\mathrm{logit}_{N}$ denotes the logit distribution obtained from the native context, and $\mathrm{logit}_{\mathrm{CD}}$ represents the final distribution used for next-token prediction.
3.2 Formal Analysis
The core component of contrastive decoding lies in the construction of the native context. Our preliminary investigation reveals that existing methods, which modify the original image, introduce several limitations (see Sec.1). Further observations and analysis show that language priors are inherent characteristics of LVLMs and remain consistent across different images, as illustrated in Fig.1. This insight leads to a natural alternative, where the native context is constructed using an image distinct from the original. Under this design, Eq. 2 is refined as:
$\mathrm{logit}_{\mathrm{CD}}(y_{t}\mid v,x,y_{<t})$ (3)
$=$ $\ (1+\alpha)\,\mathrm{logit}_{\theta}(y_{t}\mid v,x,y_{<t})$
$\quad-\alpha\,\mathrm{logit}_{\theta}(y_{t}\mid v^{\prime},x,y_{<t})$
where $v\textquoteright$ ’ is a different image from $v$. We further provide a formal analysis of the effectiveness of this approach.
Most existing methods employ sampling decoding, where a token is randomly selected according to the logit distribution. We define the set of plausible tokens under a given logit distribution as $S=V\cup T$, where $V$ represents tokens related to visual content, and $T$ represents tokens unrelated to the image, i.e., those driven by language priors. The negative context constructed by a different image yields a new logit distribution, whose plausible token set is $S\textquoteright=V\textquoteright\cup T$, given that language priors remain unchanged across images. The contrastive decoding operation can thus be interpreted as subtracting one plausible token set from another:
$S-S^{\prime}=(V\cup T)-(V^{\prime}\cup T)=V\mbox{ if }V^{\prime}\cap V=\emptyset$ (4)
In other words, this subtraction removes all language priors that are unrelated to the visual input. This approach achieves three objectives simultaneously: (1) retention of all essential visual information, (2) precise elimination of language priors, and (3) compatibility with the input distribution expected by LVLMs.
4 Method
Leveraging the cross-image consistency of language priors, we propose CICD (Cross-Images Contrastive Decoding), which effectively mitigates language priors while preventing visual information loss. In addition, we retain essential linguistic priors to ensure fluent and natural generation. As illustrated in Fig. 5, CICD comprises three key steps: (1) retrieving a distinct image, (2) identifying detrimental language priors, and (3) removing these priors through contrastive decoding.
4.1 Image Retrieval
@@
As discussed earlier, it is essential to select an image with limited information that overlaps with the input image. Since visual features are extracted by the vision encoder, a natural strategy is to measure the similarity between two images using the encoder itself to estimate their overlap. In other words, the image with the lowest similarity to the input image is selected as the reference. To ensure better generalizability, we adopt the pretrained CLIP model Radford et al. (2021) as the retriever. All images in the dataset are vectorized in advance, and during inference, the image with the lowest cosine similarity to the input is retrieved as the contrastive image.
@@
Although vector retrieval is relatively efficient given the computational cost of LVLMs, practical deployment still requires loading an additional retriever and maintaining a vector database. As illustrated in Fig.1, we observe that language priors primarily manifest in content words such as objects, attributes, and relationships. These visual elements tend to exhibit minimal overlap across natural images. Therefore, randomly selecting an image can serve as an effective and lightweight alternative for practical applications. We provide further discussion in Appendix.C.2. Note that, to ensure reproducibility, we adopt the retrieval-based strategy for consistent and deterministic image selection throughout all experiments.
4.2 Identifying Detrimental Priors
Although language priors may cause LVLMs to generate hallucinated content, they remain essential for producing fluent and coherent text. To address this, we categorize language priors into two types: essential priors and detrimental priors, with only the latter requiring elimination. Essential priors are defined as purely linguistic patterns that are entirely independent of visual input. As discussed in Sec. 1, the logit distributions associated with essential priors should remain highly consistent across diverse visual inputs. This intrinsic property provides a reliable basis for distinguishing between essential and detrimental priors.

### v5

@@ Abstract
Over-reliance on language priors is a major cause of hallucinations in Large Vision-Language Models (LVLMs), often leading to outputs that are linguistically plausible but visually inconsistent. Recent studies have explored contrastive decoding as a training-free solution. However, these methods typically construct contrastive visual inputs by perturbing the original image, resulting in distorted contrastive distributions, incomplete contrastive signals, and excessive suppression of language priors. Motivated by the observation that language priors tend to remain consistent across different images, we propose Cross-Image Contrastive Decoding (CICD), a simple yet effective training-free method that uses unrelated images as contrastive visual inputs. To address the issue of over-suppressing language priors, which can negatively affect the quality of generated responses, we further introduce a dynamic selection mechanism based on the cross-image differences in model behavior. By selectively suppressing language priors, our method reduces hallucinations without compromising the model’s performance. Extensive experiments across multiple benchmarks and LVLMs confirm the effectiveness and generalizability of CICD, particularly in image captioning, where language priors are especially dominant.  [^1] Our code is available at https://github.com/beta-nlp/CICD.
@@
Jianfei Zhao1,2, Feng Zhang1, Xin Sun1, Lingxing Kong4, Zhixing Tan4, Chong Feng1,3 1School of Computer Science and Technology, Beijing Institute of Technology 2Zhongguancun Academy 3Southeast Academy of Information Technology, Beijing Institute of Technology 4Tsinghua University {zhqingan, bit_zhangfeng, sunxin}@bit.edu.cn, klxstar@126.com, tzx.2019@tsinghua.org.cn, fengchong@bit.edu.cn
@@ 1 Introduction
Recent developments in Large Language Models (LLMs) Chiang et al. (2023a); Bai et al. (2023a) have opened a promising pathway toward Artificial General Intelligence (AGI). Building on these advances, many studies Liu et al. (2023); Dai et al. (2023); Bai et al. (2023b) have focused on integrating visual inputs into LLMs, giving rise to Large Vision-Language Models (LVLMs) capable of understanding visual content. These models can handle a wide range of vision-language tasks, including image captioning Rohrbach et al. (2018); Ye et al. (2025); Chen et al. (2025) and visual question answering (VQA) Li et al. (2023b); Lovenia et al. (2024); Yu et al. (2024).
Figure 1: Comparison of contrastive decoding–based methods. Our approach avoids overlapping visual information, training-inference gap, and excessive suppression of language priors observed in existing methods.
Despite their impressive capabilities, LVLMs are known to suffer from severe hallucinations Li et al. (2023b); Wang et al. (2023b); Leng et al. (2024). These models often generate responses based solely on memorized linguistic associations rather than actual visual input. For example, when describing an office scene, an LVLM may erroneously mention a “computer” even if no such object is present in the image. Therefore, improving the reliability and applicability of LVLMs across domains requires addressing this hallucination problem. While the causes of hallucination in LVLMs are complex, prior studies Leng et al. (2024); Wang et al. (2024b); Zhu et al. (2025a); Chen et al. (2025) have identified over-reliance on language priors as one of the primary contributing factors. Recent efforts to mitigate this issue have focused on plug-and-play, training-free methods Wang et al. (2024b); Leng et al. (2024); Zhu et al. (2025a); Zhang et al. (2025), which avoid the substantial computational costs associated with training-based approaches Wu et al. (2022); Ren et al. (2023); Chen et al. (2025). These methods Leng et al. (2024); Wang et al. (2024b); Zhu et al. (2025a) commonly employ contrastive decoding Li et al. (2023a), which highlights differences in output distributions between the original and contrastive inputs.
Various strategies Leng et al. (2024); Wang et al. (2024b); Liu et al. (2024d); Favero et al. (2024); Zhang et al. (2025) have been proposed for constructing visual inputs in contrastive decoding. While previous studies have demonstrated their effectiveness, we identify three major challenges in contrastive decoding-based methods, as illustrated in Figure 1. First, distorted distributions from contrastive visual inputs. Leng et al. (2024); Wang et al. (2024b); Liu et al. (2024d); Favero et al. (2024) propose using noised images or even removing the image entirely when computing the contrastive distribution. However, these approaches introduce a training-inference gap, as such contrastive inputs are not seen during training. This may result in distorted contrastive distributions. Second, overlapping visual information between original and contrastive inputs. Leng et al. (2024); Wang et al. (2024b); Zhang et al. (2025) propose using noisy or generated images that share visual similarities with the original inputs. However, such strategies may lead to incomplete contrastive signals. Third, trade-offs in suppressing language priors. Although over-reliance on language priors can cause hallucinations, excessively suppressing them may degrade the performance of the model. Although Zhang et al. (2025) selectively remove language priors, they augment the retained language priors, which introduces bias. As a result, mitigating language priors through contrastive decoding remains an open challenge that needs further investigation.
Figure 2: Adding the same prefix when captioning a different image leads to the same hallucination, indicating that language priors are consistent across images.
In this paper, we propose Cross-Image Contrastive Decoding (CICD), a simple yet effective method to address the aforementioned challenges in contrastive decoding. We begin by observing that LVLMs often produce highly similar outputs when given the same textual prefix but entirely different images, suggesting that language priors tend to dominate regardless of visual input (see Figure 2 for an example). Based on this observation, we construct contrastive visual inputs using different images. This approach avoids the training-inference gap and minimizes overlap between the original and contrastive inputs, thereby emphasizing the visual information in the original input and alleviating over-reliance on language priors. To prevent excessive suppression of language priors during contrastive decoding, we leverage the consistency of model behavior across images to dynamically determine, at each decoding step, whether language priors should be suppressed and to what extent. When the divergence falls below a predefined threshold, our method backs off to regular decoding. This adaptive mechanism allows CICD to effectively reduce over-reliance on language priors while preserving the model’s performance.
@@
We conduct extensive experiments across multiple benchmarks and LVLMs. The results demonstrate that our method effectively alleviates language priors, particularly in the image captioning task.
@@ 2.1 Hallucination in LVLMs
In the multimodal domain, the study of hallucinations primarily focuses on cross-modal consistency Li et al. (2023b); Wang et al. (2023b), which requires that content generated by LVLMs remains faithful to the visual input. Multiple factors contribute to hallucinations in LVLMs, including biased visual attention Zhu et al. (2025b); Li et al. (2025), attention misallocation Huang et al. (2024); Yin et al. (2025); Kang et al. (2025), and over-reliance on language priors Leng et al. (2024); Zhu et al. (2025a); Wang et al. (2024b). Our work primarily focuses on mitigating the over-reliance on language priors in LVLMs.
2.2 Language Priors in LVLMs
As LVLMs are built upon LLMs, they inherently inherit the linguistic knowledge embedded in LLMs—commonly referred to as language priors Wu et al. (2022); Ren et al. (2023). Over-reliance on these priors can cause LVLMs to generate content that is linguistically plausible but inconsistent with visual evidence. To mitigate this issue, Liu et al. (2024a); Gunjal et al. (2024); Chen et al. (2025) design specialized training datasets. However, these training-based approaches are often resource-intensive and not easily transferable to broader application scenarios. Recent studies have explored contrastive decoding Li et al. (2023a); Chuang et al. (2023) as a training-free alternative to reduce dependence on language priors. These methods construct an alternative logit distribution alongside the original one, and then contrast the two distributions to suppress language priors during decoding. The contrastive distribution is typically generated using techniques such as masking the image Leng et al. (2024), perturbing the instruction Wang et al. (2024b), augmenting the visual input Zhu et al. (2025a), or performing cross-modal conversion Zhang et al. (2025). Our approach also builds on the contrastive decoding framework. However, unlike the aforementioned methods, we propose constructing contrastive inputs using a different image and introducing a dynamic mechanism to prevent the over-suppression of language priors. This design avoids both the introduction of overlapping visual information and the training-inference discrepancies in the contrastive input, making it more effective at addressing hallucinations in LVLMs.
3 Cross-Image Contrastive Decoding
We propose Cross-Image Contrastive Decoding (CICD) for effective mitigation of language priors. CICD first retrieves a distinct image as the contrastive visual input. Then, at each decoding step, it dynamically chooses between regular decoding and contrastive decoding based on the divergence between the original and contrastive distributions. An overview of CICD is presented in Figure 3. We begin with an analysis of language priors in Section 3.1. Based on this analysis, we present methods for constructing contrastive visual inputs in Section 3.2. Finally, we discuss the importance of balanced language prior suppression in Section 3.3.
@@ 3 Cross-Image Contrastive Decoding
Figure 3: Overview of our method. We input the target image and a distinct image into the LVLM separately. During step-by-step response generation, we compute the Jensen–Shannon (JS) divergence between the two logit distributions at each step to decide whether to suppress language priors. Contrastive decoding is applied to suppress language priors, while regular decoding is used to avoid excessive suppression.
3.1 An Analysis of Language Priors
Figure 4: Similarity of the output logit distributions of LLaVA-1.5-7B in image-relevant and image-irrelevant contexts across 500 images.
While over-reliance on language priors may lead to hallucinations in LVLMs, excessive suppression of these priors can negatively affect the quality of generated texts. To investigate the effect of language priors on the generation of image-relevant and image-irrelevant tokens, we examine the logit distributions of LVLMs in both image-relevant and image-irrelevant contexts.
To this end, we randomly selected 500 images from the MSCOCO dataset and compared the similarity of the output logit distributions produced by LLaVA-1.5-7B in image-relevant and image-irrelevant contexts. We define the generation for the first token as the image-irrelevant context, and use the prefix “The picture depicts a” as the image-relevant context. We use cosine distance, Euclidean distance, and variation to study the similarity between logit distributions in the image-irrelevant and image-relevant contexts, respectively. It is evident that the logit distributions in image-irrelevant contexts exhibit high similarity across different images, regardless of the similarity metric used. In contrast, the logit distributions in image-relevant contexts do not display such consistency. We further validate this observation using additional context prompts, and the results remain consistent, suggesting that cross-image similarity can be leveraged to discriminate between image-irrelevant and image-relevant contexts. Only language priors in image-relevant contexts need to be removed.
3.2 Construction of Contrastive Visual Inputs
Based on the above analysis, we propose constructing contrastive visual inputs using an image that is distinct from the original input image. By choosing an image from a given dataset instead of synthesizing one, we avoid introducing a potential training–inference gap while maximizing visual contrast during contrastive decoding.
Intuitively, the image with the lowest similarity to the input image is preferred as the contrastive visual input for contrastive decoding. We propose two different methods for selecting a distinct image from a given dataset $\mathcal{D}_{\textrm{img}}$:
Retrieval.
@@ 3.2 Construction of Contrastive Visual Inputs
Since visual features are extracted by the vision encoder, a natural strategy is to measure the similarity between two images using the encoder itself to estimate their overlap. To ensure better generalizability, we adopt the pretrained CLIP model Radford et al. (2021) as the retriever. As all images in the dataset $\mathcal{D}_{\mathrm{img}}$ can be vectorized in advance before the inference stage, the retrieval process incurs negligible computational cost. We simply retrieve the image with the lowest cosine similarity to the input, which is then used as the contrastive image during decoding.
Random.
@@
Although vector retrieval is relatively efficient compared to the computational cost of LVLMs, practical deployment still requires loading an additional retriever and maintaining a vector database. Previous works Li et al. (2023b); Wang et al. (2023a); Ye et al. (2025) reveal that language priors primarily exist in content words such as objects, attributes, and relationships. These visual elements tend to exhibit minimal overlap across natural images. Therefore, randomly selecting an image can serve as an effective and lightweight alternative for practical applications.
3.3 Dynamic Language Prior Suppression
Although language priors may cause LVLMs to generate hallucinated content, they are essential for the model’s generation capabilities. As previously analyzed, language priors exhibit high consistency across different images in image-irrelevant contexts. Therefore, we employ the Jensen–Shannon (JS) divergence to measure the consistency between two logit distributions produced from different images $\bm{v}$ and $\bm{v}^{\prime}$, which is formally defined as
$d(\bm{v},\bm{v}^{\prime})=\mathrm{JSD}(p_{\bm{\theta}}(y_{t}|\bm{v},\mathbf{x},\mathbf{y}_{<t})\ \|\ p_{\bm{\theta}}(y_{t}|\bm{v}^{\prime},\mathbf{x},\mathbf{y}_{<t}),$ (1)
where $\mathrm{JSD}(P\|Q)=\frac{1}{2}\mathrm{KL}(P\|M)+\frac{1}{2}\mathrm{KL}(Q\|M)$, $M=\frac{1}{2}(P+Q)$, $\mathrm{KL}(P\|Q)$ denotes the KL divergence between two probability distributions $P$ and $Q$, and $p_{\bm{\theta}}(y_{t}\mid\bm{v},\mathbf{x},\mathbf{y}_{<t})$ is the probability distribution of the next token $y_{t}$ given visual input $\bm{v}$, textual input $\mathbf{x}$, and the previously generated token sequence $\mathbf{y}_{<t}$. A higher value of $d(\bm{v},\bm{v}^{\prime})$ indicates that the current context is more image-relevant, while a lower value may suggest an image-irrelevant generation.
Figure 5: JS divergence between the logit distributions of two distinct images.